# Mapping Bias in Vision Language Models: Signposts, Pitfalls, and the Road Ahead

### v1

@@ Abstract
As Vision Language Models (VLMs) gain widespread use, their fairness remains under-explored. In this paper, we analyze demographic biases across five models and six datasets. We find that portrait datasets like UTKFace and CelebA are the best tools for bias detection, finding gaps in performance and fairness between LLaVa and CLIP models. However, scene based datasets like PATA, VLStereoSet fail to be useful benchmarks for bias due to their construction. As for pronoun based datasets like VisoGender, we receive mixed signals as only some subsets of the data are useful in providing insights. To alleviate this problem, we introduce a more difficult version of VisoGender to serve as a more rigorous evaluation. Based on these results, we call for more effective and carefully designed datasets to ensure VLMs are both fair and reliable. [^1] Code for analysis and evaluation can be found at https://github.com/KuleenS/VLMBiasEval
Mapping Bias in Vision Language Models: Signposts, Pitfalls, and the Road Ahead
Kuleen Sasse1,2, Shan Chen3,4,5, Jackson Pond3,4, Danielle S. Bitterman3,4,5*thanks: Co-Senior Authors: Danielle S. Bitterman, John Osborne, John Osborne211footnotemark: 1 1Johns Hopkins University, 2University of Alabama at Birmingham, 3Harvard Medical School, 4Mass General Brigham, 5Boston Children’s Hospital Correspondence: ksasse@uab.edu
@@ 1 Introduction
In this work, we conduct a comprehensive analysis of bias in VLMs, evaluating both the models and the datasets used to assess those biases. Our contributions are threefold:
  • We analyze the biases present in five prominent VLMs across six datasets, providing new insights into model fairness.
  • We highlight the limitations of current bias evaluation datasets, demonstrating that many fail to capture the full scope of bias in multimodal models.
  • We propose future directions for developing more comprehensive and effective bias evaluation methods for VLMs, focusing on improving dataset design to mitigate conflicts between visual and textual inputs.
With this work, we aim to improve the understanding of bias in VLMs and to provide a foundation for more robust and fair models in multimodal tasks.
Figure 1: VLMs Bias Benchmark Taxonomy
@@ 2 Related Work
Bias in LLMs has been extensively studied, especially as these models have scaled in both capability and scope. However, as VLMs continue to evolve, their corresponding benchmarks have largely followed similar developmental frameworks as those used for LLMs. In this section, we review key studies on bias in LLMs and VLMs and categorize existing evaluation benchmarks based on biases related to portraits, scenes, and pronouns in VLMs.
@@ 2.2 Bias in VLMs
VLMs, which integrate visual and textual data, present a uniquely complex landscape for bias evaluation. Unlike purely text-based models, VLMs such as CLIP and DALL-E are susceptible to biases inherent in both the images and the text used during training. This dual-source bias amplifies and complicates fairness concerns, as prejudices from both modalities can influence the model’s behavior. Several studies have explored how different pretraining methods and data mixtures impact VLM performance and bias propagation in downstream tasks Wang et al. (2022b); Fabbrizzi et al. (2023); Cabello et al. (2023). These studies consistently highlight how VLMs can reinforce societal stereotypes related to gender, race, and professional roles.
For example, Wang et al. (2022a) introduced the OccupationBias dataset to assess how VLMs perpetuate occupational stereotypes. Their findings show that VLMs often associate certain professions with specific genders or races, independent of contextual information. Similarly, Fraser and Kiritchenko (2024) introduced the PAIRS dataset, which presents parallel images of individuals from different racial and gender backgrounds to evaluate how VLMs adjust outputs based on visual cues. Their work demonstrates that models frequently modify predictions and descriptions based on superficial attributes like race or gender, even when the scene remains unchanged.
Another notable dataset in this area is VisoGender Hall et al. (2023), which examines gender bias in pronoun resolution. VisoGender tests whether models correctly resolve pronouns in images of people performing professional roles, revealing significant gender biases in multimodal models. For instance, when describing a doctor and a patient, models often default to gendered pronouns based on societal stereotypes rather than visual evidence. These findings underscore the broader challenge of mitigating bias in multimodal models, where both textual and visual information contribute to predictions.
2.3 Categorization of VLM Bias Benchmarks
As shown in Figure 1, existing benchmarks for VLMs can be categorized based on the types of biases they target: portrait-based, scene-based, and pronoun-based biases. Recent research has predominantly focused on portrait-based biases in VLMs Wu et al. (2024); Han et al. (2024). Our work contributes to these efforts by examining multiple benchmarks within an organized taxonomy and identifying limitations that may lead to inaccurate measurements of a VLM’s true bias risk.
Portrait-based biases address biases related to individual appearances, such as facial features, race, gender, and age. Datasets like CelebA Liu et al. (2015) and UTKFace Zhang et al. (2017) are commonly used to evaluate these biases. CelebA, a large-scale portrait dataset, includes attributes such as "blond hair" and "heavy makeup" and is widely used to study gender bias in facial recognition models. UTKFace includes labels for age, race, and gender and is used to evaluate how consistently models identify these attributes across demographic groups.
Scene-based biases evaluate how VLMs interpret complex environments involving multiple objects or people, and how bias influences their understanding of these scenes. Datasets like VLStereoSet Tian et al. (2022) and PATA Singh et al. (2022) challenge models with images accompanied by stereotypical and anti-stereotypical captions. VLStereoSet, for example, tests whether models align stereotypical captions with particular demographics, revealing how biases manifest in everyday visual scenes. PATA extends this by including both positive and offensive captions, investigating whether models generate biased outputs based on protected categories like race or gender.
Pronoun resolution-based biases focus on gender bias in resolving pronouns, particularly in professional or occupational contexts. VisoGender is a prominent benchmark in this category Hall et al. (2023), designed to test gender bias when resolving pronouns in images of people performing professional roles. For instance, VisoGender presents images of doctors or engineers, requiring models to select the correct pronouns without relying on stereotypical gender associations. An extension, VisoGender Adversarial, introduces misleading gender labels to test model robustness against biased assumptions, forcing the model to rely solely on visual evidence rather than on stereotypical gender roles.

### v2

# debiaSAE: Benchmarking and Mitigating Vision-Language Model Bias

@@ Abstract
As Vision Language Models (VLMs) gain widespread use, their fairness remains under-explored. In this paper, we analyze demographic biases across five models and six datasets. We find that portrait datasets like UTKFace and CelebA are the best tools for bias detection, finding gaps in performance and fairness for both LLaVa and CLIP models. Scene-based datasets like PATA and VLStereoSet fail to be useful benchmarks for bias due to their text prompts allowing the model to guess the answer without a picture. As for pronoun-based datasets like VisoGender, we receive mixed signals as only some subsets of the data are useful in providing insights. To alleviate these two problems, we introduce a more rigorous evaluation dataset and a debiasing method based on Sparse Autoencoders to help reduce bias in models. We find that our data set generates more meaningful errors than the previous data sets. Furthermore, our debiasing method improves fairness, gaining 5-15 points in performance over the baseline. This study displays the problems with the current benchmarks for measuring demographic bias in Vision Language Models and introduces both a more effective dataset for measuring bias and a novel and interpretable debiasing method based on Sparse Autoencoders.  [^1] Code for analysis and evaluation can be found at https://github.com/KuleenS/VLMBiasEval
@@ 1 Introduction
In this work, we conduct a comprehensive analysis of demographic biases in VLMs, evaluating 5 models and 6 datasets used to assess those biases. We highlight the limitations of current bias evaluation datasets, demonstrating that many fail to capture the full scope of demographic bias in multimodal models. We propose future directions for developing more comprehensive and effective bias evaluation methods for VLMs, focusing on improving dataset design to mitigate conflicts between visual and textual inputs. Finally, we introduce a novel debiasing method using Sparse autoencoders (SAE) that improves fairness on two datasets across two different models.
Figure 1: VLMs Demographic Bias Benchmark Taxonomy
@@ 2 Related Work
Bias in LLMs has been studied more than in VLMs, especially as these models have scaled in both capability and scope. However, as VLMs continue to evolve, their corresponding benchmarks have largely followed similar developmental frameworks as those used for LLMs. In this section, we review key studies on bias in LLMs and VLMs and categorize existing evaluation benchmarks based on biases related to portraits, scenes, and pronouns in VLMs.
3 Related Work
Bias in LLMs has been studied more than in VLMs, especially as these models have scaled in both capability and scope. However, as VLMs continue to evolve, their corresponding benchmarks have largely followed similar developmental frameworks as those used for LLMs. In this section, we review key studies on bias in LLMs and VLMs and categorize existing evaluation benchmarks based on biases related to portraits, scenes, and pronouns in VLMs.
@@ 3.2 Bias in VLMs
VLMs, which integrate visual and textual data, present a uniquely complex landscape for bias evaluation. Unlike purely text-based models, VLMs such as CLIP and DALL-E are susceptible to learning biases from both the images and the text used during training. This dual-source complicates and potentially amplifies fairness concerns, as prejudices from both modalities can influence the model’s behavior. Several studies have explored how different pretraining methods and data mixtures impact VLM performance and bias propagation in downstream tasks Wang et al. (2022b); Fabbrizzi et al. (2023); Cabello et al. (2023); Raj et al. (2024); Hamidieh et al. (2024). These studies consistently highlight how VLMs can reinforce societal stereotypes related to gender, race, and professional roles.
Many datasets have been created to measure bias in VLMs, such as OccupationBias Wang et al. (2022a), PAIRS Fraser & Kiritchenko (2024), a synthetic dataset from Sathe et al. (2024). Another notable dataset in this area is VisoGender Hall et al. (2023), which examines gender bias in pronoun resolution. VisoGender tests whether models correctly resolve pronouns in images of people performing professional roles, revealing significant gender biases in multimodal models.
Datasets can be categorized into portrait-based, scene-based, and the aforementioned pronoun resolution based VisoGender as shown in Figure 1. Portrait based datasets like CelebA Liu et al. (2015) and UTKFace Zhang et al. (2017) are the most common to measure demographic bias in VLMs Wu et al. (2024); Han et al. (2024). These datasets use portraits of peoples’ faces to evaluate biases related to individual appearances, such as facial features, race, gender, and age. As for the scene-based datasets like VLStereoSet Tian et al. (2022) and PATA Singh et al. (2022), these datasets evaluate how VLMs interpret complex environments involving multiple objects or people, and how bias influences their understanding of these scenes by challenging the model to choose between stereotypical and non-stereotypical captions.
3.3 Sparse Autoencoders: Steering and Transfer
Sparse autoencoders have risen in popularity for use in interpreting the hidden states of large language models Bricken et al. (2023); Cunningham et al. (2023). Sparse autoencoders are shallow multilayer perceptrons with an encoder and decoder that project the hidden states of LLMs to a higher dimension sparse vector. Each dimension in this sparse vector is related to an interpretable description of what types of tokens activate that dimension. Because of their interpretable activations, they could be used for steering language models by intervening on those activations. Many different tasks have been shown to be effective for SAE steering, including code generation Karvonen et al. (2024), chatbots McGrath et al. (2024), freeform generation goals Chalnev et al. (2024), and refusal to answer certain unsafe prompts O’Brien et al. (2024).
There is very little work evaluating whether Sparse Autoencoders to can transfer to VLMs . Early work focused on transferring Sparse Autoencoders between base and chat versions of different models Kutsyk et al. (2024); Kissane et al. (2024). The only work that has successfully transferred Sparse Autoencoders to VLMs has been Gallifant et al. (2025) which has transferred Sparse Autoencoders in the context of creating classifiers for CIFAR-100 Krizhevsky (2009). However, no prior works have been investigated to debias VLMs.
@@ 4 Methods
4.1 Datasets
We consider five different image datasets that include protected attributes like sex, race, religion, and age to evaluate the fairness of these models: CelebA, PATA, UTKFace, Visogender, and VLStereoSet. Appendix A.1 has descriptions of the datasets with preprocessing steps.
4.2 Models
We evaluate the performance and biases of various VLMs across different sizes and architectures, including both open and closed-source models. For the open-source models, we assess the LLaVa v1.6 series and the PaliGemma-2 series, specifically: llava-v1.6-34b-hf, llava-v1.6-vicuna-7b, llava-v1.6-mistral-7b-hf Liu et al. (2023), google/paligemma2-10b-pt-224, and google/paligemma2-10b-pt-448 Steiner et al. (2024). Additionally, we evaluate two CLIP models, CLIP L-224 and CLIP L-336, with the latter serving as the Vision Transformer (ViT) component for all the LLaVa models mentioned. For the closed-source models, we include Google’s May 2024 release: gemini-1.5-flash-001 Radford et al. (2021); Goldin et al. (2024). For reproducibility, we set the temperature to 0 in all cases. More details about hardware and API versioning can be found in Appendix A.3
Figure 2: Diagram demonstrating our process of creating the adversarial version of the VisoGender Dataset
4.3 Prompts and Prediction
The prompts used in the experiments (shown in Table 3 in Appendix A.2) were adapted from those recommended by the LLaVa authors for model evaluation and inspiration. The initial questions for each prompt were inspired by prompts from VQA datasets like VQAv2 Goyal et al. (2017), GQA Hudson & Manning (2019), and MME Fu et al. (2024) datasets. We manually tested and refined these prompts to ensure that the models produced accurate answers on a small subset of training splits.
For multiple-choice questions(MCQ), we ask the model to choose the correct caption for the image from a list of options, ending with the recommended prompt for MCQ type questions for LLaVa: “<MCQ Options> Answer with the option’s letter from the given choices directly.”
For prediction generation, we allow models to produce a distribution over the next possible tokens. From this distribution, we select the most probable token corresponding to either (”Yes”, ”No”) for binary questions and (”A”, ”B”, ”C”, etc.) for multiple-choice questions using a process similar to the lm-evaluation-harness by EleutherAI Gao et al. (2023).
To compare our models’ effectiveness, we include a random classifier baseline. This baseline uniformly samples an answer from the label space for each question. We average across 100 runs to get an accurate estimate.
4.4 Dataset Effectiveness
Here, we define an effective dataset to measure bias as having two attributes: 1) reliance on both the visual and textual inputs and 2) difficulty.
If the model is not given the image, the prompt alone should provide little or no information that the model could make an educated guess. This fact means for effective VLM bias benchmarking datasets, models should achieve around random or lower than random accuracy. This formulation implies that for ineffective vision-language datasets, the model will be able to guess the answer from the information provided in the prompt better than random chance. This was measured by running datasets both with and without images.
Secondly, effective datasets should be challenging for the model. An easy dataset is ineffective at showing where the model fails. This leads to an underestimation of the bias in the model as any weakness is obscured by the overwhelming performance.
4.5 Adversarial Robustness
To assess model robustness, we re-evaluate performance using an adversarial version of the VisoGender dataset, as shown in Figure 2. In this experiment, we modify the prompts to intentionally mislead the model by providing incorrect gender descriptions for the individuals in the image. Specifically, we append an inaccurate gender description to the end of the question. For example, if the image shows a male doctor attending to a female patient, the modified prompt might read: ”Which one is the correct caption of this image of a woman helping a man?”
Despite this adversarial input, we expect the model to identify the correct caption, such as ”the doctor is chatting with his patient,” rather than being misled by the incorrect gender description. This evaluation forces the model to look at the image rather than relying on any text input, as the text input is misleading.
4.6 Debiasing Method
To try to remove some of the biases in the models, we propose a steering method using Sparse Autoencoders. Using the pretrained GemmaScope Sparse Autoencoders on paligemma-2-10b-224 and paligemma-2-10b-448, we hand select 12 features selecting for descriptions that mentioned either ’fairness’, ’gender equity’, or related concepts. A table of features and their descriptions selected for the models can be found in Appendix A.9.
To steer on these features, we tested five different steering methods adapted from Karvonen et al. (2024): constant, conditional per input, conditional per token, clamping, and conditional clamping . Constant adds the Sparse Autoencoder decoder associated with the chosen feature to all tokens’ hidden states scaled by a chosen coefficient. Conditional per token adds the vector from the SAE decoder to a token residual scaled by a chosen coefficient if that individual hidden state has an activation for the chosen feature beyond a limit. Conditional per input adds the vector from the SAE decoder scaled by a certain coefficient toall of the tokens’ hidden states at a layer, only if at least one of the tokens’ hidden states activate the chosen feature beyond a certain threshold. Clamping adds the decoder vector for a certain feature similar to constant, but scales the vector based on the sum of the activation of the feature and a chosen coefficient. Conditional clamping performing clamping when they exceed a certain limit on a feature. For each steering method, we test at different coefficients ranging from 4040-40- 40 to $4040$.
We focus on the VisoGender and Adversarial VisoGender datasets to benchmark our method. We evaluate the same as for the other datasets.
4.7 Metrics
4.7.1 Bias Dataset Effectiveness Metrics
5 Results
Figure 3: Performance of models on all the portrait datasets compared to their fairness (DPR demographic parity ratio, higher the better.). The dashed diagonal line represents the dividing line between trading off performance for fairness or vice versa. Arrows indicate changes in performance and fairness from CLIP to LLaVa models when the differences for both are >>> 10%. Note: PaliGemma did not use these CLIP based encoders.
5.1 Portrait Datasets
Portrait datasets exhibit key characteristics of effective datasets
Tables 4-7 in Appendix A.4 highlight two essential traits of effective datasets. First, including images leads to a substantial performance boost, with F1 scores increasing by 60-70 points, indicating that visual information is crucial for prediction. Second, these tasks are challenging for VLMs, as Macro F1 scores typically range from 60 to 80, with some reaching the 90s—suggesting significant room for improvement. Additionally, the Demographic Parity Ratio, which measures fairness, remains between 40 and 80, indicating that models are not yet perfectly fair under this metric, leaving room for further enhancements.
We are able to use these datasets to make well-founded conclusions that we discuss more in Appendix A.5. In short, adding LLMs to their base CLIP models yields more fair and more performant models. However, not all sub-tasks exhibited this improvement and therefore practitioners should be cautious when examining the overall fairness of their model using this dataset.
5.2 Scene-Based Bias Evaluation
5.2.1 VLStereoSet
As shown in Figure 4 (see Appendix), VLStereoSet is ineffective at measuring bias in VLMs). Without images, models score 50-60 on the VLBS, where lower is better, and random chance yields 33. Adding images improves the scores only slightly, reducing bias by 5-15 points. This contrasts with datasets like UTKFace and CelebA, where images boost performance by 60-70 points. This means that the text provides most of the information and, therefore, models can perform well without relying on the associated imaging data.
5.2.2 PATA
PATA may have some limitations when used as a bias evaluation dataset.
As seen in Figure 5, the PATA dataset also exhibits issues in its design, rendering it ineffective for evaluating bias in models. Across all models, performance either decreases or remains unchanged when images are included. This stagnation or drop in performance suggests that PATA does not require models to leverage the image input as intended. Moreover, the fairness evaluation (Figure 5) reveals that models tend to become less fair or show no improvement when images are provided. The fact that models can achieve nearly perfect fairness using text-only inputs suggests that the inclusion of images is irrelevant and may even confuse the models, and that this dataset is ineffective for measuring bias elicited from both text and images in VLMs.
5.3 Pronoun Resolution with VisoGender
The original VisoGender dataset is insufficient to evaluate model bias.
In Table 1, we compare the performance of our models on both the Occupation-Object and Occupation-Participant versions of the VisoGender dataset. For the Occupation-Object subset, all models, except Gemini 1.5 Flash, achieve near-perfect accuracy. The performance is similar on the Occupation-Participant dataset, with most models maintaining around 80% accuracy - except for Gemini 1.5 Flash, which performs worse than random. Gemini 1.5 Flash answers 95% of the time for the female option for both datasets, likely due to interference from its safety filtering mechanisms. Despite other datasets demonstrate bias from the VLMs, as seen in UTKFace and CelebA, VisoGender but is too easy to effectively elicit errors.
Model Dataset Acc. ↑ Avg. RB ↓ CF Acc. ↑ CF Avg. RB ↓
Gemini 1.5 Flash $O$ 49.56 -0.9209 0.00 0.0000
OP 47.33 -0.9397 21.56 -0.4266
LLaVa 34B $O$ 98.68 -0.0179 80.18 -0.4148
OP 87.76 0.0561 57.14 -0.1170
LLaVa Mistral 7B $O$ 99.12 -0.0026 92.95 -0.4198
OP 81.41 -0.2045 58.50 0.0874
LLaVa Vicuna 7B $O$ 96.48 -0.0718 62.11 -0.1051
OP 77.10 -0.3431 50.11 -0.6981
Paligemma 10B 224 $O$ 58.15 -0.7459 52.42 -0.8454
OP 61.68 -0.1361 50.11 -0.8953
Paligemma 10B 448 $O$ 54.63 -0.8907 50.66 -0.8299
OP 57.14 -0.5728 48.98 -0.9165
Random $O$ 50.00 0.0000 50.00 0.0000
OP 50.00 0.0000 50.00 0.0000
Table 1: Results for VisoGender evaluation. Acc is accuracy. CF Acc is the accuracy of the adversarial version of the dataset. The performance gap between Acc vs CF Acc shows how robust models are.
5.4 Adversarial Dataset for Pronoun Resolution
As shown in Table 1, model performance drops significantly when adversarial prompts are introduced to VisoGender, indicating an over-reliance on textual information even when it conflicts with visual cues. Most models experience a 20-30 point drop in accuracy from their baseline performance. Notably, LLaVa Mistral 7B shows the least degradation, with only a 6-point drop in accuracy on the Occupation-Object dataset, suggesting it is more robust to adversarial prompts than the other models. The significant accuracy drops across almost all models suggest that they tend to prioritize textual information over visual cues when the two are in conflict. This phenomenon, often referred to as ”sycophancy,” Sharma et al. (2023) exposes a key weakness in current bias evaluation methods: models are highly susceptible to misleading text, which skews the bias analysis. To ensure accurate assessments of bias, datasets and prompts must be meticulously designed to avoid introducing conflicting or extraneous information.
The introduction of adversarial prompts shows that most models are less robust when the visual content is at odds with the textual description, underscoring the importance of well-designed datasets for bias evaluation. The adversarial version of VisoGender provides a much-needed stress test, revealing hidden biases that are not evident in the original dataset. Across most models, we see a drop in average RB on all models. In a more difficult scenario, the models shift towards defaulting towards men as their answer. This shift highlights that these models have a stronger gender bias towards males than previously predicted by the default VisoGender dataset.
5.5 Debiasing Results
As shown in Table 2, the best debiasing method improves performance by 3-16 points on non-adversarial VisoGender and 5-12 points on adversarial versions. This highlights the promise of SAE-based steering for debiasing and robustness against adversarial attacks. Regarding fairness, RB either decreases or remains stable in sign or magnitude. Additionally, we compared SAE debiasing to debiasing using prompts by replacing standard prompts for VisoGender from 3 with those in Appendix A.10 as a baseline. Each prompt was handcrafted in an attempt to try to elicit a more fair response from the model. We evaluated each prompt on the datasets. As shown in Table 12 (Appendix A.11), even the best debiasing prompt had no effect. This indicates that SAE-based debiasing is not only more scalable and controllable, but also a higher performing pproach than prompt-based debiasing.
A scaling factor of 40 and a constant vector suffice for debiasing in many cases. The most impactful feature for debiasing related to ”terms associated with gender dynamics and equity,” aligning with the dataset’s goal of assessing gender-profession separation.
Model Intervention Steering Method Scaling Factor Feature Idx. Dataset Acc. ↑ Avg. RB ↓
Paligemma 224 None N/A N/A N/A $O$ 58.15 -0.7459
Debias Clamping -40 7624 69.16 -0.6006
None N/A N/A N/A OP 61.68 -0.1361
Debias Clamping -40 7624 64.85 -0.0092
None N/A N/A N/A Adv OO 52.42 -0.8454
Debias Constant 40 7624 67.84 0.3649
None N/A N/A N/A Adv OP 50.11 -0.8953
Debias Constant 40 7624 55.33 0.5290
Paligemma 448 None N/A N/A N/A $O$ 54.63 -0.8907
Debias Constant 40 7624 70.93 -0.0325
None N/A N/A N/A OP 57.14 -0.5728
Debias Clamping -40 7624 65.76 0.1131
None N/A N/A N/A Adv OO 50.66 -0.8299
Debias Constant 40 7624 62.11 -0.3463
None N/A N/A N/A Adv OP 48.98 -0.9165
Debias Constant 40 96024 53.51 -0.0954
Table 2: Results for debiasing evaluation on Visogender ($O$, OP) and Adversarial Visogender (Adv $O$, Adv OP). Acc is accuracy. Best debiasing method chosen for each dataset and model. Avg. RB is average resolution bias.
6 Discussion and Conclusion
This paper presents one of the first comprehensive evaluations of demographic bias in VLMs, analyzing state-of-the-art models across major datasets to assess model bias, the effectiveness of bias evaluation datasets, and promising debiasing methods.
UTKFace and CelebA datasets are effective for bias evaluation
UTKFace and CelebA effectively highlight fairness gaps between LLaVa and CLIP models, showing that VLMs are generally less biased and more performant when paired with an LLM backbone. However, biases in these models vary significantly depending on the dataset, emphasizing the need for diverse datasets to capture different bias dimensions.
VLStereoSet and PATA datasets fail to accurately measure bias
VLStereoSet and PATA suffer from data artifacts that allow models to exploit text cues, underestimating bias. Future work could refine these datasets to challenge models properly.
‘Limitations of VisoGender’
The original VisoGender dataset is no longer an effective benchmark as models are not challenged, hiding their true biases. The adversarial version of VisoGender reveals more gender-related biases, emphasizing the potential value and need for more rigorous and carefully designed datasets that challenge models to ”see” and resist being misled by textual inputs.
SAE Steering is Promising
SAE based steering on the PaliGemma-2 series of models shows promise in making models less biased and more resistant to adversarial attacks. Future work can expand on implementing this steering for more datasets and models, or implementing steering at test time based on inputs.
7 Ethics Statement
We identify the following limitations in our work. First, due to the growing popularity of multimodal models, it is impractical to evaluate all available models. We have focused on a representative subset acknowledging that this might limit the generalizability of our findings. Second, around 10 to 20 % of the images for VisoGender, PATA, and VLStereoSet were unable to be retrieved due to broken hyperlinks, possibly affecting the quality of the evaluations. Third, given the limited datasets available, we recognize that the models’ fairness discussed in this paper may not fully extend to real-world scenarios.
Appendix A Appendix
A.1 Dataset Description
A.1.1 CelebA
A.1.2 PATA
PATA is a set of about 5,000 public images crawled from the web that have been organized into 24 scenes. Each image is labeled with sex (male or female), race (White, Black, Indian, East-Asian, and Latino-Hispanic), and two age groups (youth and adults). Each image has two sets of captions associated with it: One set of positive captions with desirable descriptions of the person and one set of offensive or untoward text. These captions are developed with the protected attributes in mind and try to elicit a biased response from the model. For example, captions describing a person whose gender is the protected attribute could be a ”photo of a smart engineer”, trying to elicit a common bias that women cannot be smart and/or engineers. We use the entire dataset to run our evaluation as no training or test split is provided.
A.1.3 UTKFace
UTKFace is another portrait dataset that consists of over 20K portraits in the wild that cover a large variety of poses, facial expressions, illumination, etc. Each image is labeled with age, race, and sex. Similarly to Yang et al. (2024), we bin the age label into 5 categories: ”Child” (020)020(0-20)( 0 - 20 ), ”Young” (2040)2040(20-40)( 20 - 40 ), ”Middle-Aged” (4060)4060(40-60)( 40 - 60 ), and ”Senior” (60)limit-from60(60-)( 60 - ). We focus on classifying all three attributes and measure the bias based on the other two. We call each dataset using a two part name. The first part is what the model is predicting. The second part is what the bias is measured over. For example, ”UTKFace Race Gender” would be predicting the race of the person while measuring the difference in performance across genders. As there is no test set provided, we sample 2,000 photos of the 20,000 as a test set to run our evaluation on.
A.1.4 VisoGender
VisoGender is a set of about 700 images that are used to benchmark gender bias in VLMs. The dataset evaluates the ability of a model to correctly resolve the pronoun of a person in a picture of them working in a professional role. The dataset measures how much the model associates different occupations with different genders. They split these images into two smaller datasets: Occupation-Object and Occupation-Participant. The Occupation-Object dataset has only one person working in a profession in the image, and the model must choose that person’s pronoun. The Occupation-Participant dataset has two people in the image: one who is working in a profession and one who is participating. The task is for the model must resolve the pronoun of the one who is working. We use the entire dataset to run our prediction as no train or test split is provided.
A.1.5 VLStereoSet
A.2 Prompt Table